Find all possible jordan canonical form for the matrices whose characteristics polynomial p(t) and minimal polynomial m(t) are p(t)=(t-2)4(t-3)2 m(t)= (t-2)2(t-3)2

1
Expert's answer
2021-01-19T03:49:39-0500

First of all we remark that there is 2 eigenvalues - "Sp(A)= \{2; 3 \}". Secondly, as characteristic polynomial is of degree 6, the matrix size is "6 \times6". Let's study each of these eigenvalues separately :

  • Characteristic subspace of "\lambda=3" is of dimension 2 (by looking at the degree of "(x-3)" in the characteristic polynomial). Therefore there is only two possible Jordan forms associated to this case : "\begin{pmatrix} \lambda & 0 \\ 0 & \lambda \end{pmatrix}" or "\begin{pmatrix} \lambda & 1 \\ 0 & \lambda \end{pmatrix}". But the first case is eliminated, as that woud mean that "A|_{V_3}" ("A" restricted to the characteristic subspace associated to "\lambda=3") is diagonalizable. But as the minimal polynomial of "A" contains a term "(x-\lambda)^2", it can not be diagonalizable. Therefore the only possible Jordan form is "\begin{pmatrix} \lambda & 1 \\ 0 & \lambda \end{pmatrix} =\begin{pmatrix} 3 & 1 \\ 0 & 3 \end{pmatrix}".
  • Characteristic subspace of "\lambda=2" is of dimension 4. Therefore the possible Jordan forms are : "\begin{pmatrix} \lambda & 0 &0&0 \\ 0 & \lambda &0&0 \\0&0&\lambda&0 \\ 0&0&0&\lambda\end{pmatrix}" , "\begin{pmatrix} \lambda & 1 &0&0 \\ 0 & \lambda &0&0 \\0&0&\lambda&0 \\ 0&0&0&\lambda\end{pmatrix}" ,"\begin{pmatrix} \lambda & 1 &0&0 \\ 0 & \lambda &0&0 \\0&0&\lambda&1 \\ 0&0&0&\lambda\end{pmatrix}" , "\begin{pmatrix} \lambda & 0 &0&0 \\ 0 & \lambda &1&0 \\0&0&\lambda&1 \\ 0&0&0&\lambda\end{pmatrix}" and "\begin{pmatrix} \lambda & 1 &0&0 \\ 0 & \lambda &1&0 \\0&0&\lambda&1 \\ 0&0&0&\lambda\end{pmatrix}". The first, fourth and fifth forms are excluded by studying the minimal polynomial : first form is associated to a matrix that is diagonalizable when is restricted to "V_2" , the fourth and fifth would have the minimal polynomial containing the terms "(x-\lambda)^3" and "(x-\lambda)^4" respectively. Thus, There is only two possible Jordan forms, the second and the third.

Therefore the possible Jordan forms are :

"\begin{pmatrix} 2 & 1 &0&0 &0&0\\ 0 & 2 &0&0&0&0 \\0&0&2&1&0&0 \\ 0&0&0&2&0&0\\0&0&0&0&3&1\\0&0&0&0&0&3\end{pmatrix}" , "\begin{pmatrix} 2 & 1 &0&0 &0&0\\ 0 & 2 &0&0&0&0 \\0&0&2&0&0&0 \\ 0&0&0&2&0&0\\0&0&0&0&3&1\\0&0&0&0&0&3\end{pmatrix}" + the possible permutation of Jordan blocks.